OpenStudy (anonymous):

calculate the derivative with respect to x: y^4-2y=4x^3+x I have to use implicit differentiation

OpenStudy (anonymous):

4y^3 dy/dx - 2dy/dx = 12x^2+1 dy/dx(4y^3-2)=12x^2+1 dy/dx=(12x^2+1)/(4y^3-2) There you go!

OpenStudy (anonymous):

thanks :)